# An Overview and Discussion on Using Large Language Models for Implementation Generation of Solutions to Open-Ended Problems

### v1

@@ II Overview of Traditional Automated Implementation Generation
III Overview of Large Language Models and Diffusion Models
@@ III-A Large Language Models
In response to these challenges, a body of methods has emerged to push LLMs closer toward more sophisticated reasoning and problem-solving behaviors. This work can be broadly divided into three interrelated categories: prompting engineering, Knowledge Retrieval strategies, and Model refinement techniques (RL).
III-B Prompting Engineering
@@
Linear reasoning methods such as Chain-of-Thought (CoT) [46], Complex CoT [47], Thread-of-Thought (ThoT) [48], Chain-of-Knowledge (CoK) [49], Chain-of-Code (CoC) [50], Logical Thoughts (LoT) [51], Chain-of-Event (CoE) [52], and Chain-of-Table [53] generate a single, step-by-step sequence (chain) of responses toward the final answer. Methods differ in the type of task they target, i.e., code generation, summarization, and logical inference, and in how they refine or represent intermediate steps. CoT shows that using intermediate prompting steps can enhance accuracy, e.g., up to 39% gains in mathematical problem solving. An example in-context prompt for CoT might be: “If the problem is ‘Calculate 123 × 456,’ break it down as (100 + 20 + 3) × 456 and compute step-by-step.” Complex CoT uses more involved in-context examples, improving performance by as much as 18% on harder tasks. ThoT tackles long or chaotic contexts by breaking them into manageable parts (e.g., dividing long passages into sections for sequential summarization), while CoK strategically adapts and consolidates knowledge from multiple sources to ensure coherence and reduce hallucination. CoC specializes in code-oriented reasoning by simulating key code outputs (e.g., predicting intermediate variable states for debugging), whereas LoT integrates logical equivalences and reductio ad absurdum checks to refine reasoning chains (e.g., validating statements by identifying contradictions in their negations). CoE handles summarization by extracting, generalizing, filtering, and integrating key events (e.g., pinpointing main events from news articles), and Chain-of-Table extends CoT principles to tabular data by dynamically planning and transforming tables—such as filtering or aggregation—before generating the final answer.
@@
Branching reasoning methods, like Self-Consistency [54], Contrastive CoT (or Contrastive Self-Consistency) [55], Federated Same/Different Parameter Self-Consistency/CoT (Fed-SP/DP-SC/COT) [56], Tree-of-Thoughts (ToT) [57], and Maieutic Prompting [58], explore multiple possible reasoning paths in parallel. Branching techniques vary in how they sample or fuse paths, some relying on consensus votes and others on dynamic adaptation or tree-based elimination. Self-Consistency, for instance, samples diverse solution paths and selects the most consistent final answer, achieving gains of over 11% on math tasks. Contrastive CoT incorporates both correct and incorrect in-context examples to broaden the model’s understanding, improving performance by over 10% compared to standard CoT. Fed-SP-SC leverages paraphrased queries to crowdsource additional hints, while ToT maintains a tree of partial solutions and systematically explores them with breadth-first or depth-first strategies, offering up to 65% higher success rates than CoT on challenging math tasks. Maieutic Prompting likewise generates a tree of propositions to reconcile contradictory statements, surpassing linear methods by 20% on common-sense benchmarks.
@@
Iterative reasoning approaches, such as Plan-and-Solve (PS), Program-of-Thoughts (PoT), Chain-of-Symbol (CoS), Structured Chain-of-Thought (SCoT), and Three-Hop Reasoning (THOR), refine solutions step by step, often by passing intermediate outputs back into the model to enhance accuracy. PS explicitly decomposes tasks into planning and execution phases, where the planning phase structures the problem into smaller sub-tasks, and the execution phase solves them sequentially. This reduces semantic and calculation errors, outperforming Chain-of-Thought (CoT) prompting by up to 5% [59]. PoT enhances performance by separating reasoning from computation: the model generates programmatic solutions executed by a Python interpreter, achieving up to 12% accuracy gains in numerical and QA tasks [60]. CoS encodes spatial and symbolic relationships using concise symbolic representations, which improves reasoning in spatial tasks by up to 60.8% [61]. SCoT introduces structured reasoning through program-like branching and looping, significantly improving code generation accuracy by up to 13.79% [62]. Finally, THOR addresses emotion and sentiment analysis by splitting queries into three stages—aspect identification, opinion analysis, and polarity inference—resulting in superior performance over prior supervised and zero-shot models [63]. These approaches exemplify the power of iterative methods in breaking complex problems into manageable components, thereby reducing errors and improving overall performance.
@@
Meta-Level Guidance (MLG) methods enhance Large Language Models (LLMs) by promoting self-reflection and focusing on pertinent information, thereby reducing errors. Self-Reflection involves the model evaluating its own outputs to identify and correct mistakes, leading to improved performance. For example, in translation tasks, self-reflection enables LLMs to retrieve bilingual knowledge, facilitating the generation of higher-quality translations. Focusing is achieved through techniques like System 2 Attention (S2A)  [71], which filters out irrelevant content by prompting the model to regenerate the context to include only essential information before producing a final response. This two-step approach enhances reasoning by concentrating on relevant details, thereby improving accuracy. S2A has been shown to outperform basic prompting methods, including Chain-of-Thought (CoT) and instructed prompting, particularly on truthfulness-oriented datasets. Metacognitive Prompting (MP)  [72] introduces a five-stage process to further enhance LLM performance: (1) Comprehension: The model attempts to understands the input, ensuring clarity before proceeding; (2) Preliminary Judgment: An initial assessment is made based on the understood information; (3) Critical Evaluation: The initial judgment is scrutinized, considering alternative perspectives and potential errors; (4) Final Decision with Explanation: A conclusive decision is reached, accompanied by a rationale to support it; and (5) Self-Assessment of Confidence: The model evaluates its confidence in the final decision, reflecting on the reasoning process. This structured approach enables LLMs to perform consistently better than methods like CoT and Program Synthesis (PS) across various natural language processing tasks, including paraphrasing, natural language inference, and named entity recognition.
@@
Problem Breakdown approaches include the Least-to-Most method [73], which addresses the challenge of Chain-of-Thought (CoT) failing on problems more difficult than its exemplars by first prompting the LLM to decompose a query into sub-problems and then solving them sequentially, demonstrating notable improvements over CoT and basic prompting on tasks like commonsense reasoning and mathematical problem solving. The decompositions are characterized by their hierarchical structure, breaking down complex problems into simpler, manageable sub-tasks that build upon each other to facilitate step-by-step reasoning. Decomposed Prompting (DecomP) [74] advances this idea by delegating sub-problems to different LLMs with specialized prompts and decomposers, potentially incorporating hierarchical or recursive decomposition, or external API calls. Specialized prompts are finely tuned instructions crafted to guide each LLM toward solving specific sub-problems efficiently, focusing on unique task requirements or data contexts. Decomposers act as modular intermediaries that analyze the overall problem, partition it into distinct, logically dependent sub-tasks, and distribute them to LLMs or external tools for execution. It achieves an average 25% gain over CoT and Least-to-Most in Commonsense Reasoning. Program-Aided Language Models (PAL) [75] further leverage interleaved natural language and programmatic steps to enable Python-based execution of the reasoning process, surpassing CoT and basic methods for mathematical and commonsense tasks.
@@ III-C Knowledge Retrieval
a) Task-Specific and Schema-Based Retrieval (TSR): TSR approaches leverage structured methods to solve problems in domains such as mathematics and knowledge-intensive tasks. For instance, Schema-Based Instruction Retrieval-Augmented Generation (SBI-RAG) [80] employs schema-based instruction to solve math word problems by predicting relevant schemas, offering a structured problem-solving paradigm. Schemas, which act as templates for organizing and applying domain-specific knowledge, are inherently tied to knowledge graphs that map relationships between concepts, enhancing reasoning capabilities. The model selects the most suitable schema by aligning the problem context with predefined patterns and uses it to guide the solution process in a systematic manner. Similarly, Knowledge Graph-Enhanced RAG Framework (KRAGEN) [81] employs advanced prompting techniques, notably the graph-of-thoughts (GoT) method, to dynamically decompose complex problems into smaller subproblems. Each subproblem is addressed using relevant knowledge retrieved through the RAG framework, minimizing hallucinations and enhancing solution accuracy. The individual solutions are then consolidated to form a comprehensive answer, with KRAGEN’s graph visualization enabling users to interact with and assess the quality of the solution’s GoT structure and logic [81]. These techniques stand out for their ability to address domain-specific challenges while ensuring adaptability through schema-guided reasoning. The use of schemas not only structures the solution process but also facilitates explainability.
@@
In data-driven tasks, Generative Retrieval-Augmented Matching (GRAM) [82] addresses schema matching by employing a hierarchical classification model that dynamically generates prompts for matching attributes across schemas. Specifically, GRAM utilizes a two-step process: first, it performs a coarse-grained classification to identify potential attribute matches, and then refines these matches through fine-grained classification, enhancing the precision of schema alignment. The prompt generation is guided by large language models (LLMs), which facilitate zero-shot and few-shot learning, thereby improving efficiency and accuracy in database integration [82]. Similarly, TableRAG [83] focuses on reasoning over tabular data by retrieving and processing row-column relationships to interpret structured datasets accurately. It conducts reasoning by leveraging query expansion combined with schema and cell retrieval to pinpoint crucial information before providing it to the language models, enabling efficient data encoding and precise retrieval. This approach allows TableRAG to handle large-scale tables effectively, reducing prompt lengths and mitigating information loss during the reasoning process [83].
@@
b) Self-Aware and Adaptive Retrieval. Recent RAG frameworks emphasize self-awareness and adaptive mechanisms to address uncertainties in LLMs. Self-aware Knowledge Retrieval (SeaKR) [84] activates retrieval during high uncertainty and re-ranks snippets to ensure reliability. Specifically, SeaKR addresses uncertainties arising from the LLM’s internal state inconsistencies, triggering retrieval when the model’s self-assessed confidence is low. The re-ranking process involves selecting knowledge snippets that most effectively reduce the model’s uncertainty, thereby enhancing response accuracy [84]. Self-RAG [85] introduces iterative refinement, where retrieval queries generated during the response process enable reassessment and improvement of outputs. This reassessment involves evaluating the relevance of retrieved information during generation, allowing the model to iteratively refine its responses for enhanced accuracy. Critic-Guided Planning (CR-Planner) [86] leverages critic models to iteratively guide retrieval and reasoning toward task-specific goals. The critic model operates by evaluating potential sub-goals and their executions, assigning rewards to guide the selection of the most promising reasoning paths. This guidance ensures that the reasoning process aligns with task objectives, effectively navigating complex problem spaces [86]. For domain-specific adaptation, SimRAG [87] employs self-training, generating and filtering synthetic data to fine-tune models for specialized fields. In biomedical applications, Self-Rewarding Tree Search (SeRTS) [88] combines Monte Carlo Tree Search and Reinforcement Learning to optimize retrieval. Speculative RAG [89] improves efficiency with a two-stage process: a smaller model drafts responses, while a larger model evaluates and finalizes them. This two-step process allows the system to balance efficiency and accuracy by leveraging the strengths of both models.
@@
c) Long-Term Memory for Knowledge Retrieval: Long-term memory integration in RAG frameworks addresses the limitations of purely query-specific retrieval by enabling the retention and reuse of knowledge across tasks. HippoRAG [90], inspired by neurobiological memory structures, incorporates long-term memory directly into the retrieval process. This enables the system to consolidate and utilize past context effectively, enhancing performance in repetitive or longitudinal tasks. It transitions RAG systems from single-use retrieval mechanisms to dynamic knowledge retainers.
Various architectures embed long-term memory into RAG. MemLong [91] employs a dual-network design where a frozen LLM backbone serves as a memory encoder, while a residual side-network manages retrieval, enabling efficient caching and updating of extensive contexts (up to 65k tokens). Its key advantage is scalability without data staleness, though managing large contexts may introduce overhead. HAT [92] uses a hierarchical tree-based memory structure for recursive aggregation, enhancing coherence and summary quality through conditional traversals but may face challenges in balancing tree depth with performance. MemoRAG [93] pairs a lightweight global memory model for broad context with a powerful retrieval-generation model for focused answers, managing up to one million tokens efficiently but requiring careful system tuning. Pistis-RAG [94] emphasizes adaptability, leveraging online learning and user feedback to align responses with user preferences, though its reliance on continuous feedback may introduce variability.
@@
d) Multi-Hop and Multi-Modal Reasoning Retrieval: Multi-hop and multi-modal reasoning approaches broaden Retrieval-Augmented Generation (RAG)’s capacity to handle tasks that require complex, step-by-step deliberation and data from diverse sources. This involves performing reasoning that integrates information across multiple steps or modalities to derive comprehensive answers. Deliberation refers to the systematic process of considering various pieces of information and reasoning paths to arrive at a well-founded conclusion.
@@
Multi-layered Thoughts Enhanced RAG (METRAG) [95] integrates similarity- and utility-based reasoning for deeper contextual understanding. It does so by combining similarity-oriented retrieval with utility-oriented assessments, where a utility model, supervised by an LLM, evaluates the usefulness of retrieved documents beyond mere similarity, enhancing the relevance and quality of the information utilized in generation.
@@
RAG-Star [96] combines retrieval augmentation with Monte Carlo Tree Search (MCTS) to plan intermediate sub-queries that iteratively improve problem-solving accuracy. Accuracy, in this context, refers to the model’s ability to generate correct and relevant responses to complex queries. Retrieval augmentation involves incorporating external information into the model’s reasoning process. RAG-Star uses MCTS to explore various reasoning paths by generating and evaluating intermediate sub-queries and their potential answers, effectively guiding the model toward more accurate solutions.
@@
Knowledge Graph-Enhanced RAG Framework (KRAGEN) [81] uses Graph-of-Thoughts (GoT) methods to decompose multi-hop problems into explainable components. A GoT is a structured representation that maps out the reasoning process, storing knowledge in the form of interconnected concepts and their relationships, often derived from knowledge graphs. The GoT is constructed by the model during the reasoning process, enabling it to break down complex queries into smaller, manageable parts. This decomposition allows the model to tackle each component systematically, enhancing interpretability and the overall reasoning process.
@@
Recent research has introduced specialized frameworks that tackle the sequential nature of multi-hop queries and the integration of text and vision data. These frameworks aim to address limitations in handling complex reasoning tasks that require multiple inferential steps and the seamless combination of information from different modalities. MultiHop-RAG provides a dedicated dataset and benchmarks to rigorously assess RAG systems on multi-step queries [97], facilitating the evaluation of retrieval-augmented generation models in scenarios that necessitate reasoning across multiple documents. Retrieval-Augmented Multi-modal Chain-of-Thoughts Reasoning [98] extends Chain-of-Thought (CoT) approaches to handle images and text in tandem, enabling models to process and reason over visual and textual data simultaneously. For purely textual multi-hop question answering, HOP, UNION, GENERATE (HUG) [99] offers a three-step method that models rationales as sets of sentences, enhancing explainability without requiring explicit rationale supervision. In this framework, ”Hop” involves selecting relevant sentences, ”Union” aggregates these sentences into a coherent rationale set, and ”Generate” produces the final answer based on the aggregated rationale. The rationales modeled are the sets of sentences that collectively support the answer, providing transparency in the reasoning process by explicitly outlining the evidence considered. Multimodal-CoT and Multi-Chain Reasoning (MCR) [100] further advance reasoning by respectively separating rationale generation from answer inference for science question answering, and by prompting large language models to examine multiple parallel chains of thought before synthesizing final solutions. These approaches address complex reasoning types that require integrating diverse information sources and evaluating multiple reasoning pathways. The rationale generated includes intermediate reasoning steps that elucidate the thought process leading to the answer. Prompting is generated by designing specific instructions that guide the model to consider various perspectives and reasoning chains, thereby enhancing the robustness and accuracy of the final output.
@@
e) Self-Reflection Methods: Recent advancements underscore the value of Large Language Models (LLMs) engaging in reflective reasoning before generating a final answer. Reflective reasoning involves the model’s introspection and evaluation of its own thought processes to enhance decision-making and output quality.

### v2

@@ II Overview of Traditional Automated Implementation Generation
III Overview of Large Language Models, Prompt Engineering, Retrieval-Augmented Generation, and Reinforcement Learning
@@ III-A Large Language Models
In response to these challenges, a body of methods has emerged to push LLMs closer toward more sophisticated reasoning and problem-solving behaviors. This work can be broadly divided into three interrelated categories: Prompt Engineering, knowledge retrieval through Retrieval-Augmented Generation, and model refinement through Reinforcement Learning).
III-B Prompt Engineering
@@
Linear reasoning methods such as Chain-of-Thought (CoT), Complex CoT, Thread-of-Thought (ThoT), Chain-of-Knowledge (CoK), Chain-of-Code (CoC), Logical Thoughts (LoT), Chain-of-Event (CoE), and Chain-of-Table generate a single, step-by-step sequence (chain) of responses toward the final answer. Methods differ in the type of task they target, i.e., code generation, summarization, and logical inference, and in how they refine or represent intermediate steps. CoT shows that using intermediate prompting steps can enhance accuracy, e.g., up to 39% gains in mathematical problem solving [46]. An example of in-context prompt for CoT might be: “If the problem is ‘Calculate 123 × 456,’ break it down as (100 + 20 + 3) × 456 and compute step-by-step.” Complex CoT uses more involved in-context examples, improving performance by as much as 18% on harder tasks [47]. ThoT tackles long or chaotic contexts by breaking them into manageable parts (e.g., dividing long passages into sections for sequential summarization) [48], while CoK strategically adapts and consolidates knowledge from multiple sources to ensure coherence and reduce hallucination [49]. CoC specializes in code-oriented reasoning by simulating key code outputs (e.g., predicting intermediate variable states for debugging) [50], whereas LoT integrates logical equivalences and reductio ad absurdum checks to refine reasoning chains (e.g., validating statements by identifying contradictions in their negations) [51]. CoE handles summarization by extracting, generalizing, filtering, and integrating key events (e.g., pinpointing main events from news articles) [52], and Chain-of-Table extends CoT techniques to tabular data, dynamically applying transformations like filtering or aggregation to generate coherent answers [53].
@@
Branching reasoning methods, like Self-Consistency, Contrastive CoT (or Contrastive Self-Consistency), Federated Same/Different Parameter Self-Consistency/CoT (Fed-SP/DP-SC/COT), Tree-of-Thoughts, and Maieutic Prompting, explore multiple possible reasoning paths in parallel. Branching techniques vary in how they sample or fuse paths, some relying on consensus votes and others on dynamic adaptation or tree-based elimination. Self-Consistency, for instance, samples diverse solution paths and selects the most consistent final answer, achieving gains of over 11% on math tasks [54]. Contrastive CoT incorporates both correct and incorrect in-context examples to broaden the model’s understanding, improving performance by over 10% compared to standard CoT [55]. Fed-SP-SC leverages paraphrased queries to crowdsource additional hints [56], while ToT maintains a tree of partial solutions and systematically explores them with breadth-first or depth-first strategies, offering up to 65% higher success rates than CoT on challenging math tasks(ToT) [57]. Maieutic Prompting likewise generates a tree of propositions to reconcile contradictory statements, surpassing linear methods by 20% on common-sense benchmarks [58].
@@
Iterative reasoning approaches, such as Plan-and-Solve (PS), Program-of-Thoughts (PoT), Chain-of-Symbol (CoS), Structured Chain-of-Thought (SCoT), and Three-Hop Reasoning (THOR), refine solutions step by step, often by passing intermediate outputs back into the model to enhance accuracy. PS explicitly decomposes tasks into planning and execution phases, where the planning phase structures the problem into smaller sub-tasks, and the execution phase solves them sequentially. This reduces semantic and calculation errors, outperforming Chain-of-Thought (CoT) prompting by up to 5% [59]. PoT enhances performance by separating reasoning from computation: the model generates programmatic solutions executed by a Python interpreter, achieving up to 12% accuracy gains in numerical and QA tasks [60]. CoS encodes spatial and symbolic relationships using concise symbolic representations, which improves reasoning in spatial tasks by up to 60.8% [61]. SCoT introduces structured reasoning through program-like branching and looping, significantly improving code generation accuracy by up to 13.79% [62]. Finally, THOR tackles emotion and sentiment analysis through a three-stage approach: aspect identification, opinion analysis, and polarity inference. This structured method achieves superior performance compared to previous supervised and zero-shot models [63]. These approaches exemplify the power of iterative methods in breaking complex problems into manageable components, thereby reducing errors and improving overall performance.
@@
Meta-Level Guidance (MLG) methods enhance LLMs by promoting self-reflection and focusing on pertinent information, thereby reducing errors. Self-Reflection involves the model evaluating its own outputs to identify and correct mistakes, leading to improved performance. For example, in translation tasks, self-reflection enables LLMs to retrieve bilingual knowledge, facilitating the generation of higher-quality translations. Focusing is achieved through techniques like System 2 Attention (S2A)  [71], which filters out irrelevant content by prompting the model to regenerate the context to include only essential information before producing a final response. This two-step approach enhances reasoning by concentrating on relevant details, thereby improving accuracy. S2A has been shown to outperform basic prompting methods, including Chain-of-Thought (CoT) and instructed prompting, particularly on truthfulness-oriented datasets. Metacognitive Prompting (MP)  [72] introduces a five-stage process to further enhance LLM performance: (1) Comprehension: The model attempts to understands the input, ensuring clarity before proceeding; (2) Preliminary Judgment: An initial assessment is made based on the understood information; (3) Critical Evaluation: The initial judgment is scrutinized, considering alternative perspectives and potential errors; (4) Final Decision with Explanation: A conclusive decision is reached, accompanied by a rationale to support it; and (5) Self-Assessment of Confidence: The model evaluates its confidence in the final decision, reflecting on the reasoning process. This structured approach enables LLMs to perform consistently better than methods like CoT and Program Synthesis (PS) across various natural language processing tasks, including paraphrasing, natural language inference, and named entity recognition.
@@
Problem Breakdown approaches include the Least-to-Most method [73], which addresses the challenge of Chain-of-Thought (CoT) failing on problems more difficult than its exemplars by first prompting the LLM to decompose a query into sub-problems and then solving them sequentially, demonstrating notable improvements over CoT and basic prompting on tasks like commonsense reasoning and mathematical problem solving. The decompositions are characterized by their hierarchical structure, breaking down complex problems into simpler, manageable sub-tasks that build upon each other to facilitate step-by-step reasoning. Decomposed Prompting (DecomP) breaks complex tasks into simpler sub-tasks, each handled with tailored prompts or external tools, ensuring efficient and accurate execution. For instance, the task ”Concatenate the first letters of words in ’Jack Ryan’” is decomposed into extracting words, finding their first letters, and concatenating them [74]. DecomP leverages modular decomposers to partition problems hierarchically or recursively, assigning sub-tasks to specialized LLMs or APIs. This approach achieves a 25% improvement over CoT and Least-to-Most methods in Commonsense Reasoning. Program-Aided Language Models (PAL) [75] further leverage interleaved natural language and programmatic steps to enable Python-based execution of the reasoning process, surpassing CoT and basic methods for mathematical and commonsense tasks.
@@ III-C Retrieval-Augmented Generation
a) Task-Specific and Schema-Based Retrieval (TSR): TSR approaches leverage structured methods to solve problems in domains such as mathematics and knowledge-intensive tasks. Schema-Based Instruction Retrieval-Augmented Generation (SBI-RAG) [80] employs schema-based instruction to solve math word problems by predicting relevant schemas, offering a structured problem-solving paradigm. For instance, given the problem, “If a worker earns \$20 per hour, how much will they earn in 10 hours?”, the model predicts the multiplicative schema, which involves calculating the product of hourly earnings and hours worked. Using this schema, the problem-solving process is guided step-by-step: the model multiplies the hourly rate (\$20) by the number of hours (10), resulting in a total earning of \$200. Schemas act as templates for organizing and applying domain-specific knowledge and are inherently tied to knowledge graphs that map relationships between concepts, further enhancing reasoning capabilities. By aligning the problem context with predefined patterns, SBI-RAG ensures systematic and accurate solutions while improving explainability. Similarly, Knowledge Graph-Enhanced RAG Framework (KRAGEN) [81] employs advanced prompting techniques, notably the graph-of-thoughts (GoT) method, to dynamically decompose complex problems into smaller subproblems. Each subproblem is addressed using relevant knowledge retrieved through the RAG framework, minimizing hallucinations and enhancing solution accuracy. The individual solutions are then consolidated to form a comprehensive answer, with KRAGEN’s graph visualization enabling users to interact with and assess the quality of the solution’s GoT structure and logic [81]. These techniques stand out for their ability to address domain-specific challenges while ensuring adaptability through schema-guided reasoning. The use of schemas not only structures the solution process but also facilitates explainability.
@@
In data-driven tasks, Generative Retrieval-Augmented Matching (GRAM) addresses schema matching by employing a hierarchical classification model that dynamically generates prompts for matching attributes across schemas. Specifically, GRAM utilizes a two-step process: first, it performs a coarse-grained classification to identify potential attribute matches. For instance, given two schemas, GRAM might preliminarily match the attribute “Customer Name” in one schema with “Client Name” in another. Then, it refines these matches through fine-grained classification, analyzing the context and patterns in the data to confirm the match and enhance the precision of schema alignment. In this case, GRAM would validate that “Customer Name” and “Client Name” indeed refer to the same entity by assessing their usage and data properties. The prompt generation process, guided by LLMs, enables zero-shot and few-shot learning, allowing GRAM to perform efficiently and accurately in database integration tasks, even when minimal labeled data is available [82]. Similarly, TableRAG [83] focuses on reasoning over tabular data by retrieving and processing row-column relationships to interpret structured datasets accurately. It conducts reasoning by leveraging query expansion combined with schema and cell retrieval to pinpoint crucial information before providing it to the language models, enabling efficient data encoding and precise retrieval. This approach allows TableRAG to handle large-scale tables effectively, reducing prompt lengths and mitigating information loss during the reasoning process [83].
@@
b) Self-Aware and Adaptive Retrieval: Recent RAG frameworks emphasize self-awareness and adaptive mechanisms to address uncertainties in LLMs. Self-aware Knowledge Retrieval (SeaKR) [84] activates retrieval during high uncertainty and re-ranks snippets to ensure reliability. Specifically, SeaKR [84] addresses uncertainties arising from the LLM’s internal state inconsistencies, triggering retrieval when the model’s self-assessed confidence is low. The re-ranking process involves selecting knowledge snippets that most effectively reduce the model’s uncertainty, thereby enhancing response accuracy [84]. Self-RAG [85] introduces iterative refinement, where retrieval queries generated during the response process enable reassessment and improvement of outputs. This reassessment involves evaluating the relevance of retrieved information during generation, allowing the model to iteratively refine its responses for enhanced accuracy. Critic-Guided Planning (CR-Planner) [86] leverages critic models to iteratively guide retrieval and reasoning toward task-specific goals. The critic model operates by evaluating potential sub-goals and their executions, assigning rewards to guide the selection of the most promising reasoning paths. This guidance ensures that the reasoning process aligns with task objectives, effectively navigating complex problem spaces [86]. For domain-specific adaptation, SimRAG [87] employs self-training, generating and filtering synthetic data to fine-tune models for specialized fields. In biomedical applications, Self-Rewarding Tree Search (SeRTS) [88] combines Monte Carlo Tree Search and Reinforcement Learning to optimize retrieval. Speculative RAG [89] improves efficiency with a two-stage process: a smaller model drafts responses, while a larger model evaluates and finalizes them. This two-step process allows the system to balance efficiency and accuracy by leveraging the strengths of both models.
@@
c) Long-Term Memory for Knowledge Retrieval: Long-term memory integration in RAG frameworks addresses the limitations of purely query-specific retrieval by enabling the retention and reuse of knowledge across tasks. HippoRAG [90], inspired by the hippocampal indexing theory, integrates long-term memory by linking a knowledge graph to an LLM and prioritizing relevant nodes using the Personalized PageRank algorithm. This approach allows the model to retrieve interconnected knowledge dynamically, consolidating past context for tasks like multi-hop reasoning, achieving up to 20% better performance. It excels in repetitive and longitudinal tasks by enabling adaptive and context-aware retrieval, mimicking how the brain organizes and recalls episodic memories.
Various architectures embed long-term memory into RAG. MemLong [91] employs a dual-network design where a frozen LLM backbone serves as a memory encoder, while a residual side-network manages retrieval, enabling efficient caching and updating of extensive contexts (up to 65k tokens). Its key advantage is scalability without data staleness, though managing large contexts may introduce overhead. HAT [92] introduces a Hierarchical Aggregate Tree structure that organizes dialogue history into a tree, where each node represents aggregated information from its child nodes. This design allows the system to manage extensive conversational contexts by traversing the tree to retrieve relevant information, enhancing coherence and summary quality. The recursive aggregation enables the model to handle long-term dependencies effectively, though challenges may arise in balancing tree depth with performance. MemoRAG [93] combines a lightweight global memory model with a retrieval-generation module, using draft answers as “clues” to guide precise retrieval from extensive datasets. It efficiently handles up to one million tokens by separating memory updates from retrieval operations, ensuring scalability and contextual relevance. This architecture excels in complex tasks but requires fine-tuning to balance memory efficiency and retrieval accuracy. Pistis-RAG [94] introduces a scalable, multi-stage framework for retrieval-augmented generation (RAG) systems, emphasizing alignment with human preferences through online learning and user feedback. Its architecture comprises distinct stages—matching, pre-ranking, ranking, reasoning, and aggregating—each refining the retrieval process to enhance response quality. A notable innovation is the ranking stage, which considers both semantic relevance and LLM preferences, addressing the sensitivity of LLMs to prompt ordering. By adopting a content-centric approach, Pistis-RAG integrates user feedback to continuously adapt and align with evolving user needs, resulting in a 9.3% performance improvement on the MMLU (English) benchmark. However, the reliance on continuous user feedback may introduce variability, necessitating careful system tuning to maintain consistent performance.
@@
d) Multi-Hop and Multi-Modal Reasoning Retrieval: Multi-hop and multi-modal reasoning broaden Retrieval-Augmented Generation (RAG)’s capacity to tackle tasks requiring complex, step-by-step deliberation and integration of diverse data sources. Multi-hop reasoning connects information across multiple steps to derive coherent answers, while multi-modal reasoning combines data from various formats such as text, images, and audio. This systematic approach enhances RAG’s ability to deliver comprehensive and well-founded responses to multifaceted queries.
@@
Multi-layered Thoughts Enhanced RAG (METRAG) [95] integrates similarity- and utility-based reasoning for deeper contextual understanding. It does so by combining similarity-oriented retrieval with utility-oriented assessments, where a utility model, supervised by an LLM, evaluates the usefulness of retrieved documents beyond mere similarity using metrics like task relevance, informativeness, query-specific novelty, and completeness, enhancing the relevance and quality of the information utilized in generation.
@@
RAG-Star [96] integrates retrieval augmentation with Monte Carlo Tree Search (MCTS) to improve problem-solving accuracy by iteratively planning intermediate sub-queries. Retrieval augmentation enables the model to incorporate external information, enhancing its reasoning process. Using MCTS, RAG-Star systematically explores reasoning paths by generating and evaluating intermediate sub-queries and their potential answers. This approach balances exploration and exploitation to identify the most promising reasoning trajectories, guiding the model toward highly accurate and contextually relevant solutions.
@@
Knowledge Graph-Enhanced RAG Framework (KRAGEN) [81] employs a “Graph-of-Thoughts” methodology to decompose multi-hop reasoning problems into explainable and systematic components. This approach structures the reasoning process by representing knowledge as interconnected concepts and relationships, often derived from knowledge graphs. During problem-solving, the model constructs this graph dynamically, allowing it to break down complex queries into smaller, manageable sub-tasks. By systematically addressing each component, KRAGEN enhances both the interpretability and accuracy of its reasoning process, providing a more transparent and effective framework for handling intricate queries.
However, a potential limitation of KRAGEN is its reliance on the quality and comprehensiveness of the underlying knowledge graph. If the knowledge graph lacks certain information or contains inaccuracies, the model’s reasoning and outputs may be adversely affected. Ensuring the knowledge graph is up-to-date and accurately reflects the domain is crucial for maintaining the effectiveness of the KRAGEN framework.
Building upon the foundational concepts of multi-hop and multi-modal reasoning, recent research has proposed innovative frameworks to address the inherent complexities of such tasks. These advancements focus on refining the step-by-step reasoning process and integrating diverse modalities, enabling models to navigate intricate queries with enhanced accuracy and explainability. By tackling challenges in sequential inferencing and combining textual with visual or other modal data, these frameworks set a new benchmark for retrieval-augmented generation systems
@@
For instance, MultiHop-RAG provides a dedicated dataset and benchmarks to rigorously assess RAG systems on multi-step queries [97], facilitating the evaluation of retrieval-augmented generation models in scenarios that necessitate reasoning across multiple documents. Retrieval-Augmented Multi-modal Chain-of-Thoughts Reasoning [98] extends Chain-of-Thought (CoT) approaches to handle images and text in tandem, enabling models to process and reason over visual and textual data simultaneously. For purely textual multi-hop question answering, HOP, UNION, GENERATE (HUG) [99] offers a three-step method that models rationales as sets of sentences, enhancing explainability without requiring explicit rationale supervision. In this framework, ”Hop” involves selecting relevant sentences, ”Union” aggregates these sentences into a coherent rationale set, and ”Generate” produces the final answer based on the aggregated rationale. The rationales modeled are the sets of sentences that collectively support the answer, providing transparency in the reasoning process by explicitly outlining the evidence considered. Multimodal-CoT and Multi-Chain Reasoning (MCR) [100] further advance reasoning by respectively separating rationale generation from answer inference for science question answering, and by prompting LLMs to examine multiple parallel chains of thought before synthesizing final solutions. These approaches address complex reasoning types that require integrating diverse information sources and evaluating multiple reasoning pathways. The rationale generated includes intermediate reasoning steps that elucidate the thought process leading to the answer. Prompting is generated by designing specific instructions that guide the model to consider various perspectives and reasoning chains, thereby enhancing the robustness and accuracy of the final output.
@@
e) Self-Reflection Methods: Recent advancements underscore the value of LLMs engaging in reflective reasoning before generating a final answer. Reflective reasoning involves the model’s introspection and evaluation of its own thought processes to enhance decision-making and output quality.